# AstroClearNet: Deep image prior for multi-frame astronomical image restoration

### v1

@@ 4 Results and discussion
Such limitations suggest several paths for improving on the preliminary results. For instance, it may be possible to enhance the latent images by performing multi-frame image restoration in a higher spatial resolution, as shown in [21, 10]. This would entail modifications of AstroClearNet’s network architecture to allow for learning a super-resolved latent image $x$ with sub-pixel detail in galaxies and stars, which may facilitate the detection of small, faint sources. Moreover, an interesting avenue for future work involves leveraging the differentiability of the network (with respect to the exposures), i.e., the existence of $\partial_{y}F_{\theta}(y)$, in order to propagate uncertainties from the input exposures through the restoration process. This would allow us to construct confidence intervals for the restored latent images, providing valuable information for subsequent scientific analyses. Furthermore, as part of future work, the authors plan on rigorously comparing AstroClearNet with the ImageMM method [21], as well as exploring the possibility of combining the two methods to leverage the strengths of each.
@@ Acknowledgements
T.B. gratefully acknowledges support from the National Science Foundation (Award 1909709).

### v2

@@ 4.3 Future work
$\bullet$ Uncertainty quantification: Moreover, an interesting avenue for future work involves leveraging the differentiability of the network (with respect to the exposures), i.e., the existence of $\partial_{y}F_{\theta}(y)$, in order to propagate uncertainties from the input exposures through the restoration process. This would allow us to construct confidence intervals for the restored latent images, providing valuable information for subsequent scientific analyses.
$\bullet$ Comprehensive quantitative assessment and benchmarking: More broadly, a key avenue for the future development of AstroClearNet involves performing a comprehensive quantitative evaluation of its restoration performance. This could be pursued by comparing AstroClearNet’s latent images against restorations produced by other state-of-the-art methods, using standard image quality metrics, such as the PSNR and SSIM metrics. Alternatively, with an emphasis on photometric accuracy, one could compare source fluxes or other key photometric quantities measured from AstroClearNet’s outputs to those in established astronomical catalogs. Although we considered including such evaluations in the present study, we ultimately chose not to, as the current results do not yet match the fidelity of more established restoration methods, as shown in Section 4.2. Nevertheless, as AstroClearNet continues to mature, we intend to undertake rigorous benchmarking using both simulated and real datasets. In particular, the authors plan on rigorously comparing AstroClearNet with the ImageMM method [23], as well as exploring the possibility of combining the two methods to leverage the strengths of each.
$\bullet$ Broader applicability and future scientific utility: Looking ahead, another important direction involves extending the AstroClearNet framework to datasets beyond the Hyper Suprime-Cam (HSC). As a general-purpose multi-frame image restoration method, AstroClearNet’s ability to handle noisy and blurry images has the potential to benefit a wide range of astronomical imaging modalities. In particular, we aim to evaluate the method on data from upcoming deep optical surveys such as the Rubin Observatory’s Legacy Survey of Space and Time (LSST) [2], and are also interested in extending the framework to other wavelength regimes, including near-infrared and ultraviolet imaging, where noisy conditions are common and traditional coaddition may not achieve optimal signal-to-noise ratios. These investigations will help assess the generalization performance of AstroClearNet in the presence of diverse instrumental and atmospheric effects, and guide any necessary architectural or training adjustments to improve its adaptability across datasets.
More broadly, improved restorations produced by AstroClearNet could support critical downstream tasks in astronomy, such as the detection of faint sources, morphological classification of galaxies, and accurate photometric measurements—especially in challenging low signal-to-noise regimes. While the current implementation remains preliminary, these directions underscore the long-term scientific utility of unsupervised, data-driven restoration methods in astronomical analysis pipelines.
@@ Acknowledgements
T.B. gratefully acknowledges support from the National Science Foundation (Award 1909709 and Award 2206341).